SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Jan.10, 2006)

Attendees: Rainer Bartoldus, Al Eisner, Selina Li, Dong Su.

Operations

  • We have been running with quite low luminosity. The physics_low_lumi key is used when lumi gets below 3x10^33. Recall that the normal physics key will have a bhabha prescale of 9 on top of the flattened which already has an additional prescaling to take it down to 10 nb. For the low lumi key, it has a prescale factor of 3, rather 9. We try to save about 10 Hz of Bhabha's for rolling bhabha calibration which is still done at offline now.
  • See the DCZ section below for the only operation issues we had for the past week.

EMT

  • Rainer pointed us to a link in the TRG software page that contain information for production release used. It was indicated in that page that 18.6.1c is already been used for PC, ER, and SP production. This is the release that would have EMT prescale code in. This also means that we can change the EMT prescale to greater than 1 now.

DCZ

  • The new TSF firmware (version test_9) was installed to the TSF-X crate last Wednesday (Jan.4) during the ROD (Repair Opportunity Day). See HN posted. Marc will investigate further.
  • Today during the down time, Su Dong swapped the two TSF boards for X8 and X9 that showed up as a neighbour data problem since we worked on the TSF-X crate to load the new firmware last week. Took a cosmics run (61118) to confirm that things were working fine. The problem obviously disappeared when we started taking physics data again this evening.
  • About 200 L1 events showed TSF damages in run 60996. See HN posted.
  • We discussed a little bit about Trigger calibration validation. Su Dong reminded us that the DCH HV should be off during calibration.

Online

  • Fang tested the DCZ online lib in which the test stand dependency was removed at IR2 last Wednesday. She provided some plots to show that the data taken with the new lib were fine. The test was successful. The large deadtime seen when this was last tested on Dec.21, 2005 was probly due to incorrect trigger mask.

L3

  • Arik is working on using derivatives to improve the L3 track fitting. Maybe he will have something to show at the meeting next week if he doesn't have a time conflict with our meeting time.
  • Rainer will talk to Nicolas about TSF persistency for CM2 and L3 development. The decision was that the L1 data would go into the raw, so it would become default for the intermediate store when running Bogus/SimApp/Bear as a chain, and optional as part of Elf or Moose.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.